OpenStudy (anonymous):

Plss answer this !! Problem : Find the derivatives of the function using the limit of definition y=(3x+4)(5x-2)

OpenStudy (anonymous):

plss show ur solutions too

OpenStudy (anonymous):

\[ \lim_{\delta{x}\to0}\frac{(3(x+\delta{x})+4)(5(x+\delta{x})-2)-(3x+4)(5x-2)}{\delta{x}}\\ =\lim_{\delta{x}\to0}\frac{15x^2+30x\delta{x}+15\delta{x}^2+14x+14\delta{x}-8-15x^2-14x+8}{\delta{x}}\\ =\lim_{\delta{x}\to0}\frac{30x\delta{x}+15\delta{x}^2+14\delta{x}}{\delta{x}}\\ =\lim_{\delta{x}\to0}[30x+15\delta{x}+14]\\ =30x+14 \]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

how did you get the\[= 15x ^{2}+30xh+15h ^{2}+14x+14h-8-15x ^{2}-14x+8\]

OpenStudy (anonymous):

im confuse about this equation

OpenStudy (anonymous):

Distribute the binomial products.